The sum of the digits of a two digit number is 15. If the digits are reversed, the result exceeds the original number by 9.Find the original number.
I think the first equation is: x + y = 15

Answer by mananth(16949)


Let the digits in the units place be y
in tens place be x
when reversed
10y+x-9=10x+y
9y-9x=9
.
y-x=1
y+x=15
2y=16
y=8
x= 7
The number is 78